OpenStudy (anonymous):

could i please have help with this. Which two ordered pairs represent a proportional relationship? A.(3,1)and(6,3) B.(4,1)and(6,2) C. (3,1)and (6,2) D. (5,2)and (6,2)

OpenStudy (anonymous):

how r u doing so far and what can we help u with?

OpenStudy (anonymous):

I need help with the question i just posted becz i dont have a clue as to figuring out the proportional relationship in them numbers

OpenStudy (anonymous):

what about C?

OpenStudy (anonymous):

how did u figure it was c

OpenStudy (anonymous):

(3,1) and (6,2) \[\frac{ 3 }{ 1 }=3\] \[(6,2)=\frac{ 6 }{ 2 }=3\] so \[\frac{ 3 }{ 1 }=\frac{ 6 }{ 2 }\]